# The Developing Human Connectome Project: A Fast Deep Learning-based Pipeline for Neonatal Cortical Surface Reconstruction

### v1

@@ Abstract
The Developing Human Connectome Project (dHCP) aims to explore developmental patterns of the human brain during the perinatal period. An automated processing pipeline has been developed to extract high-quality cortical surfaces from structural brain magnetic resonance (MR) images for the dHCP neonatal dataset. However, the current implementation of the pipeline requires more than 6.5 hours to process a single MRI scan, making it expensive for large-scale neuroimaging studies. In this paper, we propose a fast deep learning (DL) based pipeline for dHCP neonatal cortical surface reconstruction, incorporating DL-based brain extraction, cortical surface reconstruction and spherical projection, as well as GPU-accelerated cortical surface inflation and cortical feature estimation. We introduce a multiscale deformation network to learn diffeomorphic cortical surface reconstruction end-to-end from T2-weighted brain MRI. A fast unsupervised spherical mapping approach is integrated to minimize metric distortions between cortical surfaces and projected spheres. The entire workflow of our DL-based dHCP pipeline completes within only 24 seconds on a modern GPU, which is nearly 1000 times faster than the original dHCP pipeline. Manual quality control demonstrates that for 82.5% of the test samples, our DL-based pipeline produces superior (54.2%) or equal quality (28.3%) cortical surfaces compared to the original dHCP pipeline.
@@ 1.2 Related Work
To overcome the limitations in the traditional neuroimage pipelines, one alternative solution is deep learning (DL), which trains a deep neural network on a large dataset and makes end-to-end predictions during inference. DL-based approaches are supported by mature frameworks such as TensorFlow [1] and PyTorch [48], and accelerated by modern graphics processing units (GPUs) with fast parallel computation. Recently, several DL-based brain MRI processing pipelines have been proposed such as FastSurfer [28], DeeepCSR [13] and iBEAT V2.0 [65]. These pipelines introduce learning-based modules to predict implicit surfaces from brain MRI for cortical surface reconstruction. The implicit surfaces are represented as a segmentation or a signed distance function (SDF), which is a level set indicating the signed distance to the surface [47]. An explicit 3D mesh can be extracted from the implicit surface using the Marching Cubes algorithm [36]. Subsequently, topology correction algorithms [57, 6, 7] are usually applied to fix topological errors in the extracted surface such that it is topologically equivalent to a sphere. However, such topology correction algorithms are usually time-consuming as they either repeatedly project the surface onto a sphere to detect topological defects [57], or iterate over all voxels to detect critical points [6, 7]. Hence, current DL-based pipelines [28, 65, 9] still require more than half an hour to process a single subject, although they can extract accurate cortical surfaces.
@@
In contrast to learning implicit surfaces, explicit learning-based approaches [67, 39, 29, 35, 54, 10, 38, 37] predict a sequence of deformations from brain MRI to deform an initial template surface explicitly into cortical surfaces. Since explicit deformations do not change the surface topology, time-consuming topology correction is not necessary if the initial surface has genus-0 topology. Therefore, explicit approaches, being both fast and capable of end-to-end learning, require only a few seconds to extract cortical surfaces, effectively circumventing the corruptions caused by imprecise segmentations. Recent explicit approaches such as CortexODE [38], CorticalFlow [35], CorticalFlow++ [54] and CoTAN [37] focus on learning diffeomorphic deformations, which have been widely used for medical image registration [8, 3, 63, 5]. Such diffeomorphic surface deformations preserve the surface topology and effectively prevent surface self-intersections, which will affect the estimation of cortical geometric features such as cortical volume, surface area and cortical thickness [38].
@@
After surface reconstruction, spherical projection or spherical mapping, is usually performed to project the cortical surface onto a spherical mesh while minimizing their metric distortions [22], e.g., the distortions of edge length, face area, or geodesic distance. The spherical projection allows the cortical surface to be represented in a spherical coordinate system for downstream tasks such as cortical surface registration and atlas construction [23, 52, 51, 11]. Conventional spherical projection approaches [22, 18] are computationally intensive due to the iterative inflation of cortical surfaces and optimization of metric distortions. Recently, [70] proposed an end-to-end unsupervised learning framework for spherical mapping. Multiscale Spherical U-Net models [71, 72] are employed to learn coarse-to-fine diffeomorphic deformation fields to deform an initial projected sphere. The Spherical U-Nets are trained by minimizing the unsupervised metric distortions between the deformed sphere and resampled cortical surfaces for each resolution.
@@ 2.1 Data Acquisition
Fig. 1: The distribution of PMA of neonates at birth and scan.
Fig. 2: The workflow of our DL-based dHCP neonatal structural MRI processing pipeline. The pipeline consists of DL-based, GPU-accelerated and CPU-based processing steps. Given a T2w structural brain MRI, the pipeline learns to extract the brain ROI and applies N4 bias field correction. The processed image is then affinely aligned to a 40-week dHCP atlas. The DL-based pipeline predicts cortical white matter and pial surfaces end-to-end from the aligned image. Next, the predicted white matter surface is iteratively inflated and smoothed to produce inflated surfaces. Based on the extracted cortical surfaces, the cortical morphological features including cortical thickness, curvature and sulcal depth are computed. The myelin map is estimated if the T1w image is provided as well. Finally, an unsupervised spherical mapping approach is adopted to project the white matter surface onto a sphere.
@@ 2.3 Dataset Preparation
We generate the initial template image, surface and sphere for training and running the pipeline. The generated templates are shared for all subjects. As depicted in Figure 3, we create an image template from a 40-week dHCP atlas [55]. To reduce the GPU memory cost, we downsample the original atlas to $0.6mm^{3}$ isotropic resolution with the size of $176\times 224\times 160$. The downsampled atlas is used as the target image template for affine registration. Next, we run the original dHCP pipeline to extract the midthickness surfaces from the dHCP atlas. For each brain hemisphere, the midthickness surface is iteratively smoothed and remeshed to create an inflated initial surface, which is in the same space as the image template. Such an initial surface mesh has 150k vertices and is used as the input template surface for cortical surface reconstruction. Finally, we use FreeSurfer mris_inflate and mris_sphere [22, 21] to project the initial surface onto an initial sphere, which has the same mesh connectivity as the initial surface with minimized metric distortions. The initial sphere is used as the input for spherical projection.
@@ 2.5.1 Diffeomorphic Surface Deformation
$\frac{\partial}{\partial t}\phi_{t}=v(\phi_{t}),~{}\phi_{0}=id,$ (1)
where $v:\Omega\rightarrow\mathbb{R}^{d}$ is a stationary velocity field (SVF), $\phi_{t}:\Omega\rightarrow\Omega$ is a deformation field for $t\in[0,T]$, and $\phi_{0}=id$ is an identity map. If the SVF $v$ is sufficiently smooth, then $\{\phi_{t}\}_{t\in{[0,T]}}$ represents an one-parameter subgroup of diffeomorphisms generated by $v$ and satisfies the group law $\phi_{s+t}=\phi_{s}\circ\phi_{t}$ [3, 5]. Therefore, we can solve the ODE (1), of which the solution is a diffeomorphism $\phi_{T}$, using the scaling and squaring method [2, 3, 5] to iteratively compute $\phi_{t}=\phi_{t/2}\circ\phi_{t/2}$. For surface deformation, the deformation field $\phi_{T}$ evolves an initial surface $S_{0}\subset\mathbb{R}^{3}$ to a target surface $S_{T}=\phi_{T}(S_{0})$ while preserving the topology of the surface. Since $\phi_{T}$ is a differentiable bijection, any two distinct points on the initial surface $S_{0}$ will not be mapped to the same location, which provides theoretical guarantee for preventing surface self-intersections.
@@ 2.5.2 Multiscale Deformation Network
Since a single SVF has limited representation ability, in this work, we propose multiscale deformation networks to learn a series of diffeomorphic deformation fields $\phi^{1},...,\phi^{L}$ for cortical surface reconstruction. As shown in Figure 5, given an affinely aligned T2w image, we clip it to the size of $112\times 224\times 160$ for the left or right brain hemispheres. The multiscale deformation network predicts cortical surfaces end-to-end from the clipped brain MRI. Note that the brain segmentations are not required as the input, which would cause subsequent corruptions on the surfaces in cases where segmentations are inaccurate.
@@
We use a 3D U-Net to extract multiscale feature maps from the input clipped T2w image. Then, a sequence of multiscale volumetric SVFs $v^{1},...,v^{L}$ are predicted from the extracted feature maps. Previous work [35, 38] adopted the forward Euler method to solve the ODE (1), which is fast but may introduce large numerical errors. Instead, we use the scaling and squaring method [2] for integration, which is more accurate and widely used for medical image registration tasks [3, 5]. For each scale $l=1,...,L$, we set the initial transformation as $\phi^{l}_{T/2^{K}}=id+\frac{T}{2^{K}}v^{l}$, and then the deformation field can be updated recurrently by $\phi^{l}_{T/2^{K-1}}=\phi^{l}_{T/2^{K}}\circ\phi^{l}_{T/2^{K}}$, where $K$ is the steps of the scaling and squaring. This is equivalent to $2^{K}$ steps of forward Euler.
@@
By integration, the network predicts multi-resolution diffeomorphic deformation fields $\phi^{l}$, each of which downsamples the original volume by a factor of $2^{L-l-1}$ for $l=1,...,L-1$. This allow the network to model coarse-to-fine surface deformations. The last deformation field $\phi^{L}$ has the same size as the input MRI volume to refine surface details. To improve the smoothness of the deformation fields, we further employ Gaussian smoothing with standard deviation $\sigma=1.0$. Finally, we apply the deformation fields iteratively to an input surface $S^{0}\in\mathbb{R}^{3}$. The displacements of the surface mesh vertices are sampled by linear interpolation. The surfaces are updated by $S^{l}=\phi^{l}(S^{l-1})$ for $l=1,...,L$ as shown in Figure 6, where $S^{L}$ is the final predicted white matter or pial surface. For white matter surface extraction, we use an initial template surface described in Figure 3 as the input surface for all subjects. The predicted white matter surfaces are further refined by Taubin smoothing [60] to improve the mesh quality. For pial surface reconstruction, we use the predicted white matter surface as the input surface. Since the initial surface has genus-0 topology and the diffeomorphic deformations are topology-preserving, all reconstructed cortical surfaces have the same topology as a 2-sphere.
@@ 2.5.3 Model Training
where $\mathcal{L}_{cd}$ is the Chamfer distance [19, 67], $\mathcal{L}_{edge}$ is the edge length loss [66, 67], $\mathcal{L}_{nc}$ is the normal consistency loss [66, 67, 37], $\lambda_{edge}$ and $\lambda_{nc}$ are weights for regularization terms. The Chamfer distance $\mathcal{L}_{cd}$ is defined as the bidirectional distance between two point sets [19]:
@@
We generate pseudo GT cortical surfaces using the original dHCP pipeline, and compute the Chamfer distance between the mesh vertices of the predicted and pseudo GT surfaces. All pseudo GT surfaces are remeshed to have 150k vertices to reduce the errors introduced by point matching during the computation of the Chamfer distance. We adopt PyTorch3D [50] which provides fast computation and gradient backpropagation for the Chamfer distance.
@@
where $v_{i},v_{j}$ are two vertices connected by the edge $(i,j)$. The normal consistency loss is defined as:
@@
Here $n_{i}$ is the unit normal vector of the face $f_{i}\in F$, and $f_{i}\cap f_{j}=(i,j)$ is the edge where the adjacent faces $f_{i}$ and $f_{j}$ intersect. The normal consistency loss penalizes the dot product of the normal vectors between any two adjacent faces.
For the architecture of the multiscale deformation network, we set $L$=4 as the number of scales. The numbers of channels for hidden layers are set to $\{16,32,64,128,128\}$ for white matter surface reconstruction and $\{16,32,32,32,32\}$ for pial surface to avoid overfitting. We use the same hyperparameters for both white matter and pial surface reconstruction. The regularization weights are set to $\lambda_{edge}$=0.3 and $\lambda_{nc}$=3.0. The number of scaling and squaring steps is set to $K$=7. We use the Adam optimizer with learning rate $10^{-4}$ to train the multiscale deformation networks for 200 epochs. After training, we select the best models with minimum validation error, i.e., minimum Chamfer distances between predicted and pseudo GT cortical surfaces.

### v2

@@ Abstract
The Developing Human Connectome Project (dHCP) aims to explore developmental patterns of the human brain during the perinatal period. An automated processing pipeline has been developed to extract high-quality cortical surfaces from structural brain magnetic resonance (MR) images for the dHCP neonatal dataset. However, the current implementation of the pipeline requires more than 6.5 hours to process a single MRI scan, making it expensive for large-scale neuroimaging studies. In this paper, we propose a fast deep learning (DL) based pipeline for dHCP neonatal cortical surface reconstruction, incorporating DL-based brain extraction, cortical surface reconstruction and spherical projection, as well as GPU-accelerated cortical surface inflation and cortical feature estimation. We introduce a multiscale deformation network to learn diffeomorphic cortical surface reconstruction end-to-end from T2-weighted brain MRI. A fast unsupervised spherical mapping approach is integrated to minimize metric distortions between cortical surfaces and projected spheres. The entire workflow of our DL-based dHCP pipeline completes within only 24 seconds on a modern GPU, which is nearly 1000 times faster than the original dHCP pipeline. The qualitative assessment demonstrates that for 82.5% of the test samples, the cortical surfaces reconstructed by our DL-based pipeline achieve superior (54.2%) or equal (28.3%) surface quality compared to the original dHCP pipeline.
@@ 1.2 Related Work
To overcome the limitations in the traditional neuroimage pipelines, one alternative solution is deep learning (DL), which trains a deep neural network on a large dataset and makes end-to-end predictions during inference. DL-based approaches are supported by mature frameworks such as TensorFlow [1] and PyTorch [54], and accelerated by modern graphics processing units (GPUs) with fast parallel computation. Recently, several DL-based brain MRI processing pipelines have been proposed such as FastSurfer [32], DeepCSR [14], SegRecon [28, 29], and iBEAT V2.0 [71]. These pipelines introduce learning-based modules to predict implicit surfaces from brain MRI for cortical surface reconstruction. The implicit surfaces are represented as a segmentation or a signed distance function (SDF), which is a level set indicating the signed distance to the surface [53]. An explicit 3D mesh can be extracted from the implicit surface using the Marching Cubes algorithm [42]. Subsequently, topology correction algorithms [63, 6, 7] are usually applied to fix topological errors in the extracted surface such that it is topologically equivalent to a sphere. However, such topology correction algorithms are usually time-consuming as they either repeatedly project the surface onto a sphere to detect topological defects [63], or iterate over all voxels to detect critical points [6, 7]. Hence, current DL-based pipelines [32, 71, 9] still require more than half an hour to process a single subject, although they can extract accurate cortical surfaces.
@@
In contrast to learning implicit surfaces, explicit learning-based approaches [73, 34, 40, 60, 45, 44, 43, 10, 11, 80] predict a sequence of deformations from brain MRI to deform an initial template surface explicitly into cortical surfaces. Since explicit deformations do not change the surface topology, time-consuming topology correction is not necessary if the initial surface has genus-0 topology. Therefore, explicit approaches, being both fast and capable of end-to-end learning, require only a few seconds to extract cortical surfaces, effectively circumventing the corruptions caused by imprecise segmentations. In order to learn high-quality manifold meshes [31], diffeomorphic deformations, which are widely used in medical image registration [8, 3, 69, 5], have been incorporated into recent explicit learning-based cortical surface reconstruction approaches such as CortexODE [44], CorticalFlow [40], CorticalFlow++ [60], CoTAN [43], V2C-Flow [11], and [80]. Such approaches aim to learn diffeomorphic surface deformations to preserve the surface topology and prevent surface self-intersections, which will affect the estimation of cortical geometric features such as cortical volume, surface area and cortical thickness [44].
@@
After surface reconstruction, spherical projection or spherical mapping, is usually performed to project the cortical surface onto a spherical mesh while minimizing their metric distortions [23], e.g., the distortions of edge length, face area, or geodesic distance. The spherical projection allows the cortical surface to be represented in a spherical coordinate system for downstream tasks such as cortical surface registration and atlas construction [24, 58, 57, 12]. Conventional spherical projection approaches [23, 19] are computationally intensive due to the iterative inflation of cortical surfaces and optimization of metric distortions. Recently, [77] proposed an end-to-end unsupervised learning framework for spherical mapping. Multiscale Spherical U-Net models [78, 79] are employed to learn coarse-to-fine diffeomorphic deformation fields to deform an initial projected sphere. The Spherical U-Nets are trained by minimizing the unsupervised metric distortions between the deformed spheres and resampled cortical surfaces for each resolution.
@@ 1.3 Contributions
Fig. 1: The distribution of PMA of neonates at birth and scan.
Fig. 2: The workflow of our DL-based dHCP neonatal brain MRI processing pipeline. The pipeline consists of DL-based, GPU-accelerated and CPU-based processing steps. Given a T2w structural brain MRI, the pipeline learns to extract the brain ROI (Figure 4) and applies N4 bias field correction. The processed image is then affinely aligned to a 40-week dHCP atlas. The DL-based pipeline predicts cortical white matter and pial surfaces end-to-end from the aligned image (Figure 5). Next, the white matter surface is iteratively inflated and smoothed to produce inflated surfaces. Based on predicted cortical surfaces, the cortical morphological features including cortical thickness, curvature and sulcal depth are computed. The myelin map is estimated if the T1w image is provided. Finally, an unsupervised spherical mapping approach is adopted to project the white matter surface onto a sphere (Figure 7).
@@ 2.3 Dataset Preparation
We generate the initial template image, surface and sphere for training and running the pipeline. The generated templates are shared for all subjects. As depicted in Figure 3, we create an image template from a 40-week dHCP atlas [61]. As we aim to develop a fast and memory-efficient pipeline, we downsample the original atlas from $0.5mm^{3}$ to $0.6mm^{3}$ isotropic resolution and clip it to the size of $176\times 224\times 160$ to reduce the GPU memory cost. The downsampled atlas is used as the target image template for affine registration. It should be noted that the downsampling is not necessary if GPU memory allows.
Next, we run the original dHCP pipeline to extract the midthickness surfaces from the dHCP atlas. For each brain hemisphere, the midthickness surface is iteratively smoothed and remeshed to create an inflated initial surface, which is in the same space as the image template. Such an initial surface mesh has 150k vertices and is used as the input template surface for cortical surface reconstruction. This enables the proposed DL-based pipeline to predict cortical surfaces with the same mesh connectivity and fixed number of vertices, while the original dHCP pipeline extracts cortical surfaces with adapted number of vertices, leading to inaccurate estimation of cortical morphological features such as sulcal depth for younger subjects. Finally, we use FreeSurfer mris_inflate and mris_sphere [23, 22] to project the initial surface onto an initial sphere, which has the same mesh connectivity as the initial surface with minimized metric distortions. The initial sphere is used as the input for spherical projection.
@@ 2.5.1 Diffeomorphic Surface Deformation
$\frac{\partial}{\partial t}\phi_{t}=v_{t}(\phi_{t}),~{}\phi_{0}=id,$ (1)
where $v_{t}:\Omega\times[0,T]\rightarrow\mathbb{R}^{d}$ is a time-varying velocity field, $\phi_{t}:\Omega\rightarrow\Omega$ is a deformation field for $t\in[0,T]$, and $\phi_{0}=id$ is an identity map. In this work, we only consider stationary velocity fields (SVF), i.e., $v_{t}=v:\Omega\rightarrow\mathbb{R}^{d}$. If the SVF $v$ is sufficiently smooth, then $\{\phi_{t}\}_{t\in{[0,T]}}$ represents an one-parameter subgroup of diffeomorphisms generated by $v$ and satisfies the group law $\phi_{s+t}=\phi_{s}\circ\phi_{t}$ [3, 5]. Therefore, we can solve the ODE (1), of which the solution is a diffeomorphism $\phi_{T}$, using the scaling and squaring method [33, 2, 3, 5] to recursively compute $\phi_{t}=\phi_{t/2}\circ\phi_{t/2}$. For surface deformation, the deformation field $\phi_{T}$ evolves an initial surface $S_{0}\subset\mathbb{R}^{3}$ to a target surface $S_{T}=\phi_{T}(S_{0})$ while preserving the topology of the surface. Since $\phi_{T}$ is a differentiable bijection, any two distinct points on the initial surface $S_{0}$ will not be mapped to the same location, which provides theoretical guarantee for preventing surface self-intersections.
@@ 2.5.2 Multiscale Deformation Network
Since a single SVF has limited representation ability, in this work, we propose multiscale deformation networks to learn a series of diffeomorphic deformation fields $\phi^{1},...,\phi^{L}$ for cortical surface reconstruction. As shown in Figure 5, given an affinely aligned T2w image, we clip it from the original size of $176\times 224\times 160$ to $112\times 224\times 160$ for either left or right brain hemisphere. Afterwards, the multiscale deformation network predicts cortical surfaces end-to-end from the clipped brain MRI. Since the proposed neural network model only processes one brain hemisphere at a time, the clipped brain MRI not only provides necessary information for cortical surface reconstruction, but also effectively reduces both GPU memory costs and inference time. Note that the brain segmentations are not required as the input, which would cause subsequent corruptions on the surfaces in cases where segmentations are inaccurate.
@@
We use a 3D U-Net to extract multiscale feature maps from the input clipped T2w image. Then, a sequence of multiscale volumetric SVFs $v^{1},...,v^{L}$ are predicted from the extracted feature maps. Each SVF $v^{l}$ downsamples the original brain MRI volume by a factor of $2^{L-l-1}$ for $l=1,...,L-1$. The last SVF $v^{L}$ has the same size as the input volume to capture fine details. Previous work [40, 44] adopted the forward Euler method to solve the ODE (1), which is fast but may introduce large numerical errors. Instead, we use the scaling and squaring method [33, 2] for integration, which is more accurate and widely used for medical image registration tasks [3, 5]. For each scale $l=1,...,L$, we set the initial transformation as $\phi^{l}_{T/2^{K}}=id+\frac{T}{2^{K}}v^{l}$, and then the deformation field can be updated recurrently by $\phi^{l}_{T/2^{K-1}}=\phi^{l}_{T/2^{K}}\circ\phi^{l}_{T/2^{K}}$, where $K$ is the steps of the scaling and squaring. This is equivalent to $2^{K}$ steps of forward Euler.
@@
By integration, the network predicts multi-resolution diffeomorphic deformation fields $\phi^{l}$, each of which has the same size as the corresponding SVF $v^{l}$. To improve the smoothness of the deformation fields, we further employ Gaussian smoothing with standard deviation $\sigma=1.0$. Finally, we apply the deformation fields iteratively to an input surface $S^{0}\subset\mathbb{R}^{3}$. The displacements of the surface mesh vertices are sampled by linear interpolation. The surfaces are updated by $S^{l}=\phi^{l}(S^{l-1})$ for $l=1,...,L$. We set the number of scales to $L=4$, which allows the multiscale deformation network to model coarse-to-fine surface deformations. The deformation fields $\phi^{1}$ and $\phi^{2}$ have been downsampled by a factor of 4 and 2 respectively. As shown in Figure 6, $\phi^{1}$ and $\phi^{2}$ contain global information and deform the input surface $S^{0}$ into coarse and smooth intermediate surfaces. $\phi^{3}$ and $\phi^{4}$, which have the same size as the input brain MRI, capture detailed cortical folds and provide further refinement for the final predicted white matter or pial surface $S^{L}$.
For white matter surface extraction, we use an initial template surface described in Figure 3 as the input surface for all subjects. The predicted white matter surfaces are further refined by Taubin smoothing [66] to improve the mesh quality, i.e., the smoothness and regularity of the triangular mesh. For pial surface reconstruction, we use the predicted white matter surface as the input surface. Since the initial surface has genus-0 topology and the diffeomorphic deformations are topology-preserving, all reconstructed cortical surfaces have the same topology as a 2-sphere.
@@ 2.5.3 Model Training
where $\mathcal{L}_{cd}$ is the bidirectional Chamfer distance [20, 72, 73], $\mathcal{L}_{edge}$ is the edge length loss [72, 73], $\mathcal{L}_{nc}$ is the normal consistency loss [72, 73, 43], $\lambda_{edge}$ and $\lambda_{nc}$ are weights for regularization terms. The Chamfer distance $\mathcal{L}_{cd}$ is defined as the bidirectional distance between two point sets $X,Y\subset\mathbb{R}^{3}$ [20]:
@@
Specifically, $X$ and $Y$ are the sets of mesh vertices of the predicted and pseudo GT cortical surfaces respectively. The pseudo GT cortical surfaces are generated using the original dHCP pipeline [50]. All pseudo GT surfaces are remeshed to have 150k vertices to reduce the errors introduced by point matching during the computation of the Chamfer distance. We adopt PyTorch3D [56] which provides fast computation and gradient backpropagation for the Chamfer distance.
@@
where $v_{i},v_{j}$ are two vertices connected by the edge $e_{ij}=(i,j)$. The normal consistency loss is defined as:
@@
Here $n_{i}$ is the unit normal vector of the face $f_{i}\in F$, and $e=f_{i}\cap f_{j}$ is the edge where the adjacent faces $f_{i}$ and $f_{j}$ intersect. The normal consistency loss penalizes the dot product of the normal vectors between any two adjacent faces. The value of the normal consistency loss is determined by the PMA of neonates and the curvature of cortical surfaces. For older neonates with highly folded and curved cortical surfaces, there is a large difference between the normal orientations of two adjacent faces. Consequently, the normal consistency loss reaches a high value to enforce surface smoothness. For younger subjects with immature cortex, the loss is relatively lower as the cortical surfaces are already sufficiently smooth.
For the architecture of the multiscale deformation network, we set $L=4$ as the number of scales as discussed in Section 2.5.2. A smaller $L$ would offer insufficient global information, while a larger $L$ is more suitable for brain MRI at higher image resolution. The numbers of channels for hidden layers are set to $\{16,32,64,128,128\}$ for white matter surface reconstruction and $\{16,32,32,32,32\}$ for pial surface to avoid overfitting. We use the same hyperparameters for both white matter and pial surface reconstruction. The regularization weights are selected by grid search, and set to $\lambda_{edge}=0.3$ and $\lambda_{nc}=3.0$ to achieve satisfactory mesh quality. The number of scaling and squaring steps is set to $K=7$ following [5]. Empirically [44, 43], we use the Adam optimizer with learning rate $10^{-4}$ to train the multiscale deformation networks for 200 epochs. After training, we select the best models with minimum validation error, i.e., minimum Chamfer distances between predicted and pseudo GT cortical surfaces.